a student heated 21.5 g pure hydrated calcium sulfate ,CaSO4 x H2O .17.0 g anhydrous calcium sulfate was formed .Determine the molecular formula for this hydrated salt
1
Expert's answer
2019-10-09T07:36:28-0400

CaSO4•H2O->(t) CaSO4+H2O

m(CaSO4)=17g

m(CaSO4•xH2O)=21.5g

M(CaSO4•xH2O)=m(CaSO4•xH2O)*M(CaSO4)/m(CaSO4)=21.5*136/17=172g/mol

x=(M(CaSO4•xH2O)-M(CaSO4))/18=(172-136)/18=2

CaSO4•2H2O